OpenStudy (anonymous):

How do you graph -6x+3y=-6

OpenStudy (anonymous):

Replace x with 0 to find the y intercept -6(0) + 3y = -6 3y = -6 /3 /3 y = -2 The y-intercept would be at (0,-2) Replace y with 0 to find the x intercept -6x + 3(0) = -6 -6x = -6 /-6 /-6 x = 1 The x intercept is at (0,1) Plot the y-intercept and x-intercept on a graph and draw a line through the two points and ta dum you have your graph :3

OpenStudy (anonymous):

oooh wait

OpenStudy (anonymous):

the x-intercept is at (1,0) sorry

OpenStudy (anonymous):

Thanks you sooo much